# A Bayesian latent Gaussian conditional autoregressive copula model for analyzing spatially-varying trends in rainfall

### v1

@@ 8 Concluding remarks
Evaluating the availability of rainfall water is essential for rainfed agriculture, which is predominantly practiced in India. Considering potential trends in rainfall patterns due to climate change, we develop a statistical model to analyze annual total rainfall data for 34 meteorological subdivisions of mainland India from 1951 to 2014. Our approach models the marginal distributions using a gamma regression model while capturing dependence through a Gaussian conditional autoregressive (CAR) copula model. Given the natural variation in average annual rainfall across the country’s diverse climatic regions, we incorporate areally-varying gamma regression coefficients within a latent Gaussian model framework. The dependence structure in both the likelihood and prior layers is governed by the neighborhood relationships of the regions, where we examine both CAR and intrinsic CAR structures for the priors. We employ Markov chain Monte Carlo algorithms to facilitate Bayesian inference, precisely a combination of block Metropolis-Hastings and single-component Metropolis-Hastings steps within a Gibb sampler. Additionally, the proposed methodology effectively imputes missing data within the Gibbs sampling. Our model demonstrates superior performance in simulation studies compared to alternatives that do not account for dependence structures in the data or prior layers. Applying our methodology to the Indian areal rainfall dataset, we estimate model parameters, correctly quantify the uncertainties, and investigate the potential impact of climate change on rainfall patterns across India. The mean annual rainfall shows a significant negative trend for three meteorological subdivisions of northeastern India.
@@
A drawback of the proposed methodology is a specific dependence structure across space determined by the neighborhood structure of India’s meteorological subdivisions. The literature contains a large number of copula models for Gaussian graphical models. However, our approach is parsimonious and implementable in all scenarios where a CAR model is applicable. The significance of our proposed methodology extends beyond rainfall modeling. It provides a generalizable framework for analyzing other spatially dependent climatic variables, such as temperature and humidity, which exhibit complex dependence structures. This approach is instrumental in climate impact assessments, where accurate predictions of future rainfall trends can inform adaptation and mitigation strategies [60]. Incorporating spatially varying coefficients (SVCs) into rainfall modeling is crucial for accurately capturing the inherent heterogeneity of precipitation patterns across different regions. Traditional models with fixed coefficients often fail to account for local variations, leading to oversimplified representations of rainfall processes. Allowing model parameters to change spatially enables SVC models to represent better the complex interactions between environmental factors and rainfall distribution. This enhanced flexibility improves the precision of hydrological predictions, which is vital for effective water resource management and flood risk assessment. For instance, studies have demonstrated that SVC models can effectively capture both short-range and long-range spatial dependencies in hydrological processes, leading to more accurate estimations of mean annual runoff coefficients [61, 45, 62]. Integrating SVCs into rainfall models allows for a more nuanced understanding of spatial variability, leading to more reliable and region-specific hydrological insights.
@@ Acknowledgement
The authors would like to thank Aritra Basak for providing some codes for preliminary analysis and the adjacency matrix used in our paper. The research of the second author is supported by the Start-up Research Grant (SRG), Science and Engineering Research Board, Department of Science & Technology, Government of India, with Award No. SERB-MATH-2023632.
@@
Codes (written in R) for implementing our proposed methodology and the dataset analyzed in this paper are provided in the Supplementary Material.
Disclosure statement

### v2

@@ 8 Concluding remarks
Evaluating the availability of rainfall water is essential for rainfed agriculture, which is predominantly practiced in India. Considering potential trends in rainfall patterns due to climate change, we develop a statistical model to analyze monsoon total rainfall data for 34 meteorological subdivisions of mainland India from 1951 to 2014. Our approach models the marginal distributions using a gamma regression model while capturing dependence through a Gaussian conditional autoregressive (CAR) copula model. Given the natural variation in monsoon total rainfall across the diverse climatic regions of the country, we incorporate gamma regression with spatially varying coefficients (SVCs) within a latent Gaussian model framework. The dependence structure in both the likelihood and prior layers is governed by the neighborhood relationships of the regions, where we examine both CAR and intrinsic CAR structures for the priors. We employ Markov chain Monte Carlo algorithms to facilitate Bayesian inference, precisely a combination of block Metropolis-Hastings and single-component Metropolis-Hastings steps within a Gibbs sampler. Additionally, the proposed methodology effectively imputes missing data within the Gibbs sampling. Our model demonstrates superior performance in simulation studies compared to alternatives that do not account for dependence structures in the data or prior layers. Applying our methodology to the Indian areal rainfall dataset, we estimate model parameters, correctly quantify the uncertainties, and investigate the potential impact of climate change on rainfall patterns across India. The mean annual rainfall shows a significant negative trend for three meteorological subdivisions of northeastern India.
@@
A drawback of the proposed methodology is a specific dependence structure across space determined by the neighborhood structure of India’s meteorological subdivisions. The literature contains a large number of copula models for Gaussian graphical models. However, our approach is parsimonious and implementable in all scenarios where a CAR model is applicable. The significance of our proposed methodology extends beyond rainfall modeling. It provides a generalizable framework for analyzing other spatially dependent climatic variables, such as temperature and humidity, which exhibit complex dependence structures. This approach is instrumental in climate impact assessments, where accurate predictions of future rainfall trends can inform adaptation and mitigation strategies [fischer2013robust]. Incorporating SVCs into rainfall modeling is crucial for accurately capturing the inherent heterogeneity of precipitation patterns across different regions. Traditional models with fixed coefficients often fail to account for local variations, leading to oversimplified representations of rainfall processes. Allowing model parameters to change spatially enables SVC models to better represent the complex interactions between environmental factors and rainfall distribution. This enhanced flexibility improves the precision of hydrological predictions, which is vital for effective water resource management and flood risk assessment. For instance, studies have demonstrated that SVC models can effectively capture both short-range and long-range spatial dependencies in hydrological processes, leading to more accurate estimations of mean annual runoff coefficients [bakar2020interpolation, gelfand2003spatial, arnaud2002influence]. Integrating SVCs into rainfall models enables a more nuanced understanding of spatial variability, resulting in more reliable and region-specific hydrological insights.
Our model specification incorporates spatial structure at two levels: in the marginal parameters through CAR/ICAR priors, and in the data layer through the CAR-copula dependence. This two-layer dependence is conceptually appealing, reflecting epistemic uncertainty in regional parameters and aleatoric uncertainty in rainfall totals. Similar hierarchical structures have been studied in spatial statistics [reich2006effects, hughes2013dimension, hanks2015restricted], where spatial random effects can coexist at multiple levels. While our results indicate that both layers contribute complementary information, potential confounding between them cannot be entirely ruled out. Approaches such as sum-to-zero constraints or restricted spatial regression [reich2006effects, hughes2013dimension] could help formally orthogonalize the two components, and we view this as an important direction for future work. In this work, we have focused on a gamma regression framework with time as the primary covariate, but the approach naturally extends to a broader class of models. Specifically, any distribution from the exponential family can be used for the marginal distributions, with the mean parameter linked to multiple temporal covariates (such as year, global climate indices, or atmospheric variables) through a generalized linear model structure. Spatial variation can be incorporated by modeling regression coefficients as SVCs with appropriate spatial priors. Moreover, dependence across regions can be captured through a Gaussian copula with CAR or GMRF correlation structures, combined with marginal distributions drawn from the exponential family. This generalization enables more flexible spatio-temporal modeling of rainfall or other environmental responses within a unified framework.
@@ Acknowledgement
The authors thank an Associate Editor and four anonymous Reviewers for suggesting several crucial changes that substantially improved the manuscript. Additionally, the authors would like to thank Aritra Basak for providing the relevant codes for preliminary analysis and the adjacency matrix used in our paper. The research of the second author is supported by the Start-up Research Grant (SRG), Science and Engineering Research Board, Department of Science & Technology, Government of India, with Award No. SERB-MATH-2023632.
@@
Codes (written in R) for implementing our proposed methodology and the dataset analyzed in this paper are provided in the Supplementary Material. Additionally, the extension of gamma regression to a more general framework for non-Gaussian areal datasets, further exploratory analysis, univariate trace plots with effective sample sizes, and model diagnostics are also provided.
Conflict of interest statement